OpenStudy (loser66):

Negate this sentence: "The integer n is odd only if 3n is odd" I need method to get the answer, not the answer itself Please, help

OpenStudy (loser66):

@perl

OpenStudy (anonymous):

The integer n is odd only if 3n is even ??

OpenStudy (loser66):

How to get it?

OpenStudy (perl):

p only if q means ' if p then q '

OpenStudy (loser66):

This is what I did: p: the integer n is odd q: 3n is odd p--> q then negate it

OpenStudy (anonymous):

write it as an "if ... then..." statement first, then you will unconfuse yourself

OpenStudy (anonymous):

I am not sure if it is right or not.. It is what I think..

OpenStudy (perl):

how do we negate ' if p then q ' ? the negation is p and (not q)

OpenStudy (anonymous):

the "only if" is annoying and confusing change it to "if P then Q" and you can negate it easily

OpenStudy (loser66):

p--> = ~p or q

OpenStudy (loser66):

negate the right hand side to gate p and not q

OpenStudy (loser66):

but it 's weird when verbalizing it.

OpenStudy (anonymous):

If 3n is odd, then integer n is odd...

OpenStudy (perl):

Let p = n is odd q = 3n is odd the original statement is 'if n is odd then 3n is odd. ' The negation is ' n is odd and 3n is not odd '

OpenStudy (perl):

so the negation is n is odd and 3n is even

OpenStudy (loser66):

I am not allowed to use "Not" in my sentence

OpenStudy (anonymous):

"not odd' we call "even" on planet earth

OpenStudy (perl):

the opposite of odd (for integers) is even . if n is not odd, then n is even

OpenStudy (loser66):

@perl so that what I got, but it is weird, right? how n is odd and 3n is even, hahaha??

OpenStudy (perl):

and we are assuming we are talking about ingegers

OpenStudy (perl):

loser, yes it is weird, because it is false . (the original statement is true)

OpenStudy (perl):

if you negate a true statement, you will get a false statement.

OpenStudy (perl):

this is just a logical exercise ;)

OpenStudy (perl):

if a statement is true, then its negated form will be false

OpenStudy (loser66):

ooooohhh.... yessssssssssss!

OpenStudy (loser66):

Thank you everybody. :)

OpenStudy (perl):

lets prove the original statement is true. Given integer n is odd, then by definition of odd integer n = 2k+1 for some integer k. But now 3n = 3(2k+1) And 3n =3(2k+1) = 6k + 3 = 6k + 2 + 1 = 2(3k+1) + 1 So we have 3n = 2*m + 1 , where m = 3k+1 so 3n is an odd number by definition of odd integer.

OpenStudy (perl):

its good to practice these proofs on your own

OpenStudy (perl):

and since this is a true statement (since we proved it), if you negate it you know it will be false

OpenStudy (loser66):

I have proving problems like this too. It's not too hard to go over. But this problem confused me and your statement " negate true will get a false one" enlightens me. Thanks again.